OpenStudy (anonymous):

WILL SOMEONE PLEASE HELP ME WORK THROUGH THIS PROBLEM?! Sharon measures out some aluminum pellets an records a mass of 12.8g. She then measures the volume to be 4.85 mL. Calculate the density an percent error, if the actual density of the aluminum is 2.7g/mL. Show the work used to arrive to this answer.

OpenStudy (anonymous):

Take note that density is mass over volume. Using her obtained values, her density was 12.8/4.85=2.639g/mL. Then use ratio and proportion of actual value to obtained value. So we say 2.7 as to 100% And 2.639 as to $ The $=2.639*100/2.7=97.7%. Therefore the percentage error is 100-97.7=2.3%

OpenStudy (anonymous):

thank you so much!!!!!!

OpenStudy (anonymous):

My pleasure.